the suggested dose or erythromycin is 40 mg /kg/day q6h A neonate weighing 3000 g would...

  1. the suggested dose or erythromycin is 40 mg /kg/day q6h
    1. A neonate weighing 3000 g would receive ___ mg per dose?
    2. Erythromycin is provided as 10 mg / mL
    3. How many mL would you give?
Answer #1

Suggested dose of Erythromycin=40mg/kg/day q6h

Weight of Neonate=3000g = 3kg

As 1 kg = 1000 g

Then as per weight of neonate per dose will be =40X 3= 120 mg per day q6h

Per dose amount will be=120/4= 30 mg

Erythromycin provided as= 10 mg/ml

Then for 120 mg required ml will be =30/10=3 ml per dose